OpenStudy (anonymous):

Use factoring to fnd the zeros of y=x^2+x-12 and the axis of symmetry .

OpenStudy (anonymous):

y=x^2+x-12 y=(x+4)(x-3)=0 x=-4, and x=3

OpenStudy (anonymous):

to check it you can use the quadratic formula again

OpenStudy (anonymous):

/any question lysa

OpenStudy (anonymous):

- Thee only question I have is if you can look at the rest .

OpenStudy (anonymous):

ok